OpenStudy (peace420):

Please helpp . A train travels at 100 miles per hour. An equation can be written that compares the time (t) with the distance (d). Define the domain and range. The domain is time (t) and the range is 100. The domain is 100 and the range is time (t). The domain is distance (d) and the range is time (t). The domain is time (t) and the range is distance (d).

OpenStudy (peace420):

it is not.. i tried it and it wasn't right

OpenStudy (stephcurry):

...My apologies, try the 3rd one and see...

OpenStudy (peace420):

okay hope its right

OpenStudy (peace420):

could you help with another?

OpenStudy (stephcurry):

Sure.

OpenStudy (peace420):

Write the Standard Form of the line with x-intercept of 3 and y-intercept of 4. 3x + 4y = 12 4x + 3y = 12 3x + 4y = 7 4x + 3y = 7

OpenStudy (stephcurry):

What is the y intercept ?

OpenStudy (stephcurry):

But I believe it's the first one...

OpenStudy (peace420):

i also tried that and it was wrong

OpenStudy (stephcurry):

Try the third one.

OpenStudy (peace420):

okay, one more?

OpenStudy (stephcurry):

Go ahead.

OpenStudy (peace420):

A train travels at 100 miles per hour. Write an equation that compares the time (t) with the distance (d). d = t + 100 d = 100t t = 100d t = d + 100

OpenStudy (stephcurry):

I think that's the third one too.

OpenStudy (peace420):

it is not :(

OpenStudy (stephcurry):

The second one.

OpenStudy (peace420):

My first try i only got 1 right now i only got 2 right

OpenStudy (stephcurry):

That is good progress.

OpenStudy (peace420):

i need 100% not 50's :(

OpenStudy (stephcurry):

I have to go, if you don't mind giving me a medal, I would really appreciate it. Good luck on your next questions.:)